OpenStudy (anonymous):

fin a1 in geometric series. s12=3071.25 and r=2 a.0.5 b.0.75 c.2.5 d.3

OpenStudy (anonymous):

\[S _{n}=\frac{ a(r^n-1) }{ r-1 }\] \[S _{12}=\frac{ a(2^{12}-1) }{ 2-1 }\] \[3071.25=\frac{ a(2^{12}-1) }{ 2-1 }\] Solve for a by rearranging the expression.